My APR is 20%, compounded daily. What is my annual effective interest rate (APY or ia)?
Answer #1

ia = (1 + r / m)m - 1

   = (1 + 0.20 / 365)365 - 1

   = 1.2213 - 1

   = 0.2213   or 22.13%